I cant get the last two questions for this one.

For the given pipe assembly, all coordinates are in feet and the magnitude of the applied couple moment around pipe CD is MCDM_{CD}M​CD​​. Calculate:

NE z = F_Ř lb C(-2,4, 2H) k Falb Mcp lb. ft A(0, 0, H) D (-3,5, 1.5H) F2 lb B(0, 3, H) 0 (0, 0, 0) F = -Fk lb

F1= 50lb F2= 70lb MD=150lb*ft H=5

1. The resultant moment vector for the couple moments acting on the pipe assembly

2. The resultant moment vector for the system at O
Answer #1

M Fi x PERPENDI CULAR DSIAN CE (235x sok 100502 1 F2xPERreNDi cULAR D)STANCE 703 E2+3 =-140 k- 3502 =-350- 140 150 JOME, M DR